# A mathematical theory of topological invariants of quantum spin systems

### v2

@@ 6.4 Topological invariants of gapped states on subsets of ${\mathbb{R}}^{n}$
So far we assumed that $\Lambda$ is an arbitrary countable subset of ${\mathbb{R}}^{n}$ with a uniform $O(r^{n})$ bound on the number of points in a ball of radius $r$. Suppose $\Lambda\subset W^{\epsilon}$ for some CS set $W$ and some $\epsilon>0$. If this is the case, we will say the pair $(\Lambda,\{V_{j}\}_{j\in\Lambda})$ describes a quantum lattice system on $W$. Then for any ${\mathsf{F}}\in{\mathfrak{D}}_{al}$ the component ${\mathsf{F}}^{X}$ vanishes for any brick $X$ which does not intersect $W^{\epsilon}\cap\Lambda$, and thus for any $U\in{\mathcal{C}\mathcal{S}}_{n}$ we have ${\mathfrak{D}}_{al}(U)={\mathfrak{D}}_{al}(U\wedge W)$ and ${\mathfrak{D}}^{\psi}_{al}(U)={\mathfrak{D}}^{\psi}_{al}(U\wedge W)$. Thus the assignment $U\mapsto{\mathfrak{D}}^{\psi}_{al}(U)$ can be regarded as a local Lie algebra over the site ${{{\mathcal{C}\mathcal{S}}_{n}}/W}$.
In particular, if a compact Lie group $G$ acts locally on the quantum lattice system on $W$ and preserves a gapped state $\psi$, the generator of the action ${\mathsf{Q}}:{\mathfrak{g}}\rightarrow{\mathfrak{D}}^{\psi}_{al}$ takes values in the sub-algebra ${\mathfrak{D}}^{\psi}_{al}(W)$. Consequently ${\mathfrak{D}}^{\psi,\bf G}_{al}$ is a graded local Fréchet-Lie algebra over ${{{\mathcal{C}\mathcal{S}}_{n}}/W}$ pointed by ${\mathsf{Q}}\in{\mathfrak{D}}^{\psi,\bf G}_{al}(W)$. Picking a cover ${\mathfrak{U}}$ of $W$ and applying the Čech functor gives an acyclic pointed DGLA whose commutator class can be evaluated on any CS cocycle $\beta$ of $\hat{W}$ to give an inner derivation. Its $\psi$-average is a $G$-invariant polynomial on ${\mathfrak{g}}$ which depends only on the class of $({\mathfrak{U}},\beta)$ in $H^{\bullet}(\hat{W},{\mathbb{R}})$. Thus a quantum lattice system on $W$ has a topological invariant which is a linear function of $H^{\bullet}(\hat{W},{\mathbb{R}})\rightarrow{\rm Sym}^{\bullet}{\mathfrak{g}}^{*}[2]$. It is easy to check that this linear function has degree $3$.
For example, consider a $U(1)$-invariant gapped state of a quantum lattice system on $W={\mathbb{R}}^{2}$ affinely embedded in ${\mathbb{R}}^{3}$. The topological invariant defined in Section 6.2 is obtained by contracting with a 2-cocycle of $S^{2}$ (the sphere at infinity) and vanishes for dimensional reasons. On the other hand, by contracting with the 1-cocycle of $\hat{W}=S^{1}$ one obtains the Hall conductance of this system.
For a more non-trivial example, for any $n>2$ consider a finite graph $\Gamma\subset S^{n-1}$ whose edges are geodesics and take $W$ to be the cone with base $\Gamma$ and apex at an arbitrary point of ${\mathbb{R}}^{n}$. The invariants of $U(1)$-invariant gapped states of quantum lattice systems on $W\subset{\mathbb{R}}^{n}$ are labeled by generators of the free abelian group $H^{1}(\Gamma,{\mathbb{Z}})$. This example goes beyond TQFT since $W$ need not be smooth or even locally Euclidean.
@@
First let us characterize ${\mathfrak{D}}_{al}(U)$ in terms of 0-chains. A 0-chain on ${\mathbb{Z}}^{n}$ is an element ${\mathsf{a}}=\{{\mathsf{a}}_{j}\}_{j\in{\mathbb{Z}}^{n}}\in\prod_{j\in{\mathbb{Z}}^{n}}{\mathfrak{D}}_{al}(\{j\})$ such that
$\|{\mathsf{a}}\|_{k}:=\sup_{j\in{\mathbb{Z}}^{n}}\|{\mathsf{a}}_{j}\|_{\{j\},k}<\infty.$ (25)
We say a 0-chain ${\mathsf{a}}$ is supported on $U$ if ${\mathsf{a}}_{i}=0$ whenever $i\notin U$, and write $C_{0}(U)$ for the set of $U$-supported 0-chains, endowed with the norms (25) for $k\geq 0$.
Proposition A.1.
Let $U\subset{\mathbb{R}}^{n}$ be nonempty and let $U^{1}:=\{x\in{\mathbb{R}}^{n}:d(x,U)\leq 1\}$ be its 1-thickening.
@@
    $(\partial{\mathsf{f}})^{Y}:=\sum_{j\in{\mathbb{Z}}^{n}\cap U}{\mathsf{f}}_{j}^{Y}$
@@
Lemma A.1.
For any nonempty $U,Y\subset{\mathbb{R}}^{n}$ we have
$1+d(Y,U^{1}\cap{\mathbb{Z}}^{n})+\operatorname{diam}(Y)\leq 2(1+d(Y,U)+\operatorname{diam}(Y))$
@@
Choose $y\in Y$ and $u\in U$ with $d(y,u)=d(Y,U)$, and choose $z\in{\mathbb{Z}}^{n}$ with $d(u,z)\leq 1$. Then since $z\in U^{1}\cap{\mathbb{Z}}^{n}$ we have
@@
Then either ${\mathsf{f}}_{i}^{Y}=0$ or $d(Y,U^{1}\cap{\mathbb{Z}}^{n})=d(Y,j)$ and $\|{\mathsf{f}}_{i}^{Y}\|=\|{\mathsf{F}}^{Y}\|$, and so
$\|{\mathsf{f}}_{i}\|_{k}$ $=\sup_{Y\in{\mathbb{B}}_{n}\backslash\{\varnothing\}}(1+\operatorname{diam}(Y)+d(Y,\{i\})^{k}\|{\mathsf{f}}_{i}^{Y}\|$
$=\sup_{Y\in{\mathbb{B}}_{n}\backslash\{\varnothing\}}(1+\operatorname{diam}(Y)+d(Y,U^{1}\cap{\mathbb{Z}}^{1}))^{k}\|{\mathsf{F}}^{Y}\|,$
which by Lemma A.1 is bounded by $2^{k}\|{\mathsf{F}}\|_{U,k}$.
$ii)$. Suppose that ${\mathsf{f}}$ is a $U$-supported 00-chain. For any $k\geq 0$ we have
$\|\partial{\mathsf{f}}^{Y}\|$ $\leq\sum_{j\in{\mathbb{Z}}^{n}\cap U}\|{\mathsf{f}}_{j}^{Y}\|$
$\leq\|{\mathsf{f}}\|_{k+2n+1}\sum_{j\in{\mathbb{Z}}^{n}\cap U}(1+\operatorname{diam}(Y)+d(j,Y))^{-k-2n-1}$
$\leq\|{\mathsf{f}}\|_{k+2n+1}(1+\operatorname{diam}(Y)+d(U,Y))^{-k}\sum_{j\in{\mathbb{Z}}^{n}\cap U}(1+\operatorname{diam}(Y)+d(j,Y))^{-2n-1}$
$\leq\|{\mathsf{f}}\|_{k+2n+1}(1+\operatorname{diam}(Y)+d(U,Y))^{-k}(1+\operatorname{diam}(Y))^{-n}\sum_{j\in{\mathbb{Z}}^{n}}(1+d(j,Y))^{-n-1}$
$\leq\|{\mathsf{f}}\|_{k+2n+1}(1+\operatorname{diam}(Y)+d(U,Y))^{-k}(1+\operatorname{diam}(Y))^{-n}\sum_{j\in{\mathbb{Z}}^{n}}(1+d(j,Y))^{-n-1}$
It is not hard to show that for any brick $Y$ the quantity $(1+\operatorname{diam}(Y))^{-n}\sum_{j\in{\mathbb{Z}}^{n}}(1+d(j,Y))^{-n-1}$ is bounded by a constant $C$ depending only on $n$, which shows $\|\partial{\mathsf{f}}^{Y}\|\leq C\|{\mathsf{f}}\|_{k+2n+1}$. ∎
Proposition A.1 will allow us to apply the results of [11] on 0-chains. The results in [11] are phrased in terms of the norms
$\|{\mathsf{a}}\|^{KS}_{x,k}:=\sup_{r>0}(1+r)^{k}\inf_{{\mathsf{b}}\in{\mathfrak{d}}(B_{r}(x))}\|{\mathsf{a}}-{\mathsf{b}}\|$
where ${\mathfrak{d}}(B_{r}(x))$ is the set of traceless anti-hermitian operators strictly localized on the ball of radius $r$ around $x\in{\mathbb{R}}^{n}$. To apply their results we prove the equivalence of these norms:
Lemma A.2.
@@
where $C,C^{\prime}$ are constants depending only on $k$.
@@
Suppose $\|{\mathsf{a}}\|_{\{x\},k+2n+1}<\infty$ and let $r>0$. Define ${\mathsf{b}}:=\sum_{X\subset B_{r}(x)}{\mathsf{a}}^{X}$. Then
$\|{\mathsf{b}}-{\mathsf{a}}\|\leq\|{\mathsf{a}}\|_{\{x\},k+2n+2}\sum_{X\nsubseteq B_{r}(x)}(1+\operatorname{diam}(X)+d(x,X))^{-k-2n-2}$
@@
$\leq C(1+r)^{-k}\|{\mathsf{a}}\|_{\{x\},k+2n+2}$
where in the last line we used Lemma 3.1. This proves (27). To prove (28) suppose $\|{\mathsf{a}}\|^{KS}_{x,k}<\infty$ and let $X$ be any brick. Set $r:=\lfloor(\operatorname{diam}(X)+d(x,X))/4\rfloor$. Then $X\nsubseteq B_{r}(x)$. Indeed, if $X\subseteq B_{r}(x)$ then $d(x,X)\leq r$ and $\operatorname{diam}(X)\leq 2r$, implying $\operatorname{diam}(X)+d(x,X)\leq 3r$, which is impossible. Choose ${\mathsf{b}}\in{\mathfrak{d}}(B_{r}(x))$ with $\|{\mathsf{a}}-{\mathsf{b}}\|\leq(1+r)^{-k}\|{\mathsf{a}}\|_{x,k}^{KS}$. Since $X\nsubseteq B_{r}(x)$ we have ${\mathsf{b}}^{X}=0$, and so
@@
$\leq 4^{n+2k}\|{\mathsf{a}}\|_{x,k}^{KS}(1+\operatorname{diam}(X)+d(x,X))^{-k},$
where in the second line we used [11, Proposition C.1]. ∎
A.2 Proof of Lemma 3.7
[…]
Proof of Lemma 3.7.
$\mathcal{J}$ and $\mathcal{K}$ correspond to $\mathscr{J}_{{\mathsf{H}},w_{\Delta}}$ and $\mathscr{J}_{{\mathsf{H}},W_{\Delta}}$ in Section 4.1 of [11]. Part $i)$ follows from the definition of $\mathcal{K}$ and the fact that ${\mathsf{H}}$ preserves $\psi$. Part $ii)$ follows from Proposition A.1 and Lemma A.2, together with [11, Lemma F.1] (specifically line (177) therein). Part $iii)$ is [11, line (72)]. ∎
@@ Appendix B Inhomogeneous Maurer-Cartan equation
Let $({\mathfrak{M}},{\mathsf{B}})$ be a pointed pronilpotent DLGA which is a limit of an inverse system of nilpotent pointed DGLAs $({\mathfrak{M}}_{N},{\mathsf{B}}_{N})$, $N\in{\mathbb{N}}$. The set ${\rm MC}({\mathfrak{M}},{\mathsf{B}})$ of ${\mathsf{B}}$-twisted MC elements has an additional equivalence relation. This section revolves around this additional structure leading eventually to a proof of Prop. 6.1.
@@
A ${\mathsf{B}}$-twisted MC element ${\mathsf{p}}$ gives rise to a degree $-1$ derivation $\partial_{\mathsf{p}}=\partial+{\rm ad}_{\mathsf{p}}$ of ${\mathfrak{M}}$ which squares to zero (twisted differential).
Lemma B.1.
@@
We are going to define an equivalence relations on ${\rm MC}({\mathfrak{M}},{\mathsf{B}})$ and ${\rm MC}({\mathfrak{M}}_{N},{\mathsf{B}}_{N})$ for all $N$. This is done in the same way as for the ordinary (homogeneous) Maurer-Cartan equation [29, 30].
@@
Second, there are Lie algebra homomorphisms from ${\mathfrak{M}}_{0}$ (resp. ${\mathfrak{M}}_{N,0}$) to the Lie algebras of affine-linear vector fields on ${\mathfrak{M}}_{-1}$ (resp. ${\mathfrak{M}}_{N,-1}$). This homomorphism maps ${\mathsf{a}}\in{\mathfrak{M}}_{0}$ or ${\mathfrak{M}}_{N,0}$ to the affine-linear vector field
$\xi_{\mathsf{a}}({\mathsf{p}})=[{\mathsf{a}},{\mathsf{p}}]-\partial{\mathsf{a}},$
@@
We say that elements $p_{1},p_{2}$ of ${\rm MC}({\mathfrak{M}},{\mathsf{B}})$ or ${\rm MC}({\mathfrak{M}}_{N},{\mathsf{B}}_{N})$ are equivalent if they belong to the same orbit of these actions.
Remark B.1.
@@
Now come the interesting statements. Assume from now on that the DGLAs ${\mathfrak{M}}_{N}$ and ${\mathfrak{M}}$ are acyclic, that ${\mathfrak{j}}_{N}\subset{\mathfrak{M}}_{N}$ is central for all $N>1$, and that the morphisms $r_{N,N-1}$ are surjective for all $N>1$.
Lemma B.4.
@@
The only if statement follows from Lemma B.1. To prove the if direction, we use induction on $N$. Assume ${\rm MC}({\mathfrak{M}}_{N-1},{\mathsf{B}}_{N-1})$ is non-empty. Let ${\mathsf{p}}_{N-1}\in{\rm MC}({\mathfrak{M}}_{N-1},{\mathsf{B}}_{N-1})$. Pick $\tilde{\mathsf{p}}_{N}\in r_{N,N-1}^{-1}({\mathsf{p}}_{N-1})$. Since ${\mathsf{p}}_{N-1}$ is a ${\mathsf{B}}_{N-1}$-twisted MC element and $r_{N,N-1}({\mathsf{B}}_{N})={\mathsf{B}}_{N-1}$, we must have
$\partial\tilde{\mathsf{p}}_{N}+\frac{1}{2}[\tilde{\mathsf{p}}_{N},\tilde{\mathsf{p}}_{N}]={\mathsf{B}}_{N}+{\mathsf{q}}_{N}$
for some ${\mathsf{q}}_{N}\in{\mathfrak{j}}_{N}$. We look for a solution of the ${\mathsf{B}}_{N}$-twisted MC equation of the form ${\mathsf{p}}_{N}=\tilde{\mathsf{p}}_{N}+{\mathsf{b}}_{N}$ where ${\mathsf{b}}_{N}\in{\mathfrak{j}}_{N}$. Taking into account that ${\mathfrak{j}}_{N}$ is central, the inhomogeneous MC equation reduces to $\partial{\mathsf{b}}_{N}=-{\mathsf{q}}_{N}$. Since by assumption ${\mathfrak{j}}_{N},{\mathfrak{M}}_{N}$, and ${\mathfrak{M}}_{N-1}$ form a short exact sequence and the latter two are acyclic, so is ${\mathfrak{j}}_{N}$. Hence such a ${\mathsf{b}}_{N}$ exists. This completes the inductive step proving that ${\rm MC}({\mathfrak{M}}_{N},{\mathsf{B}}_{N})\neq\varnothing$ for all $N$. Moreover, we also proved that the morphisms ${\rm MC}({\mathfrak{M}}_{N},{\mathsf{B}}_{N})\rightarrow{\rm MC}({\mathfrak{M}}_{N-1},{\mathsf{B}}_{N-1})$ are surjective. Therefore by Lemma B.1 ${\rm MC}({\mathfrak{M}},{\mathsf{B}})$ is non-empty. ∎
The above lemma proves the first part of Prop. 6.1. Indeed when $[{\mathfrak{M}}_{1},{\mathfrak{M}}_{1}]=0$, ${\rm MC}({\mathfrak{M}}_{1},{\mathsf{B}}_{1})\neq\varnothing$ is implied by acyclicity. Next we show that with the above assumption on ${\mathfrak{M}}_{N}$ and ${\mathfrak{M}}$ all ${\mathsf{B}}$-twisted MC elements are equivalent.
Lemma B.5.
@@
$\exp({\mathsf{b}}+{\mathsf{a}})({\mathsf{p}})=\exp({\mathsf{b}})({\mathsf{p}})-\partial{\mathsf{a}}.$

### v3

@@ 6.4 Topological invariants of gapped states on subsets of ${\mathbb{R}}^{n}$
Consider a $U(1)$-invariant gapped state of a quantum lattice system on $W={\mathbb{R}}^{2}$ affinely embedded in ${\mathbb{R}}^{3}$. If we regard it as a quantum lattice system on ${\mathbb{R}}^{3}$, then the topological invariant defined in Section 6.2 vanishes for dimensional reasons. On the other hand, by contracting with the 1-cocycle of $\widehat{W}=S^{1}$ one obtains the Hall conductance of this system.
For a more non-trivial example, for any $n>2$ consider a finite graph $\Gamma\subseteq S^{n-1}$ whose edges are geodesics and take $W$ to be the cone with base $\Gamma$ and apex at an arbitrary point of ${\mathbb{R}}^{n}$. The invariants of $U(1)$-invariant gapped states of quantum lattice systems on $W\subseteq{\mathbb{R}}^{n}$ are labeled by generators of the free abelian group $H^{1}(\Gamma,{\mathbb{Z}})$. This example goes beyond Chern-Simons field theory, since $W$ need not be smooth or even locally Euclidean.
@@
First let us characterize ${\mathfrak{D}}_{al}(U)$ in terms of 0-chains. A 0-chain on ${\mathbb{Z}}^{n}$ is an element ${\mathsf{a}}=\{{\mathsf{a}}_{x}\}_{x\in{\mathbb{Z}}^{n}}\in\prod_{x\in{\mathbb{Z}}^{n}}{\mathfrak{D}}_{al}(\{x\})$ such that
$\|{\mathsf{a}}\|_{k}:=\sup_{x\in{\mathbb{Z}}^{n}}\|{\mathsf{a}}_{x}\|_{\{x\},k}<\infty.$ (142)
We say a 0-chain $\mathsf{a}$} is supported on $U$ if ${\mathsf{a}}_{x}=0$ whenever $x\notin U$, and write $C_{0}(U)$ for the set of $U$-supported 0-chains, endowed with the norms (142) for $k\geq 0$.
Proposition 119.
Let $U\subseteq{\mathbb{R}}^{n}$ be nonempty and let $U^{1}:=\{x\in{\mathbb{R}}^{n}:d(x,U)\leq 1\}$ be its 1-thickening.
@@
    $(\partial{\mathsf{f}})^{Y}:=\sum_{x\in{\mathbb{Z}}^{n}\cap U}{\mathsf{f}}_{x}^{Y}$ (143)
@@
Lemma 120.
For any nonempty $U\subseteq{\mathbb{R}}^{n}$ and any $Y\in{\mathbb{B}}_{n}$ we have
$1+d(Y,U^{1}\cap{\mathbb{Z}}^{n})+\operatorname{diam}(Y)\leq 2(1+d(Y,U)+\operatorname{diam}(Y))$ (144)
@@
Since the distances and thickenings involving $U$ depend only on the closure of $U$, we may assume that $U$ is closed. Choose $y\in Y$ and $u\in U$ with $d(y,u)=d(Y,U)$, and choose $z\in{\mathbb{Z}}^{n}$ with $d(u,z)\leq 1$. Then since $z\in U^{1}\cap{\mathbb{Z}}^{n}$ we have
@@
Then either ${\mathsf{f}}_{i}^{Y}=0$ or $d(Y,U^{1}\cap{\mathbb{Z}}^{n})=d(Y,i)$ and $\|{\mathsf{f}}_{i}^{Y}\|=\|{\mathsf{F}}^{Y}\|$, and so
[…]
which by Lemma 120 is bounded by $2^{k}\|{\mathsf{F}}\|_{U,k}$.
ii)ii). Suppose that $\mathsf{f}$} is a $U$-supported 0-chain. For any $k\geq 0$ we have
$\|(\partial{\mathsf{f}})^{Y}\|$ $\leq\sum_{x\in{\mathbb{Z}}^{n}\cap U}\|{\mathsf{f}}_{x}^{Y}\|$
$\leq\|{\mathsf{f}}\|_{k+2n+1}\sum_{x\in{\mathbb{Z}}^{n}\cap U}(1+\operatorname{diam}(Y)+d(x,Y))^{-k-2n-1}$
$\leq\|{\mathsf{f}}\|_{k+2n+1}(1+\operatorname{diam}(Y)+d(U,Y))^{-k}\sum_{x\in{\mathbb{Z}}^{n}\cap U}(1+\operatorname{diam}(Y)+d(x,Y))^{-2n-1}$
$\leq\|{\mathsf{f}}\|_{k+2n+1}(1+\operatorname{diam}(Y)+d(U,Y))^{-k}(1+\operatorname{diam}(Y))^{-n}\sum_{x\in{\mathbb{Z}}^{n}}(1+d(x,Y))^{-n-1}.$ (148)
It is not hard to show that for any brick $Y$ the quantity $(1+\operatorname{diam}(Y))^{-n}\sum_{x\in{\mathbb{Z}}^{n}}(1+d(x,Y))^{-n-1}$ is bounded by a constant $C$ depending only on $n$, which shows $\|\partial{\mathsf{f}}\|\leq C\|{\mathsf{f}}\|_{k+2n+1}$. ∎
Proposition 119 will allow us to apply the results of [19] on 0-chains. The results in [19] are phrased in terms of the norms
$\|{\mathsf{a}}\|^{KS}_{x,k}:=\sup_{r>0}(1+r)^{k}\inf_{{\mathsf{b}}\in{{\mathfrak{d}}_{l}}(B_{r}(x))}\|{\mathsf{a}}-{\mathsf{b}}\|$ (149)
where ${{\mathfrak{d}}_{l}}(B_{r}(x))$ is the set of traceless anti-hermitian operators strictly localized on the ball of radius $r$ around $x\in{\mathbb{R}}^{n}$. To apply their results we prove the equivalence of these norms:
Lemma 121.
@@
where $C,C^{\prime}$ are constants depending only on $n$.
@@
Suppose $\|{\mathsf{a}}\|_{\{x\},k+2n+2}<\infty$ and let $r>0$. Define ${\mathsf{b}}:=\sum_{X\subseteq B_{r}(x)}{\mathsf{a}}^{X}$. Then
$\|{\mathsf{b}}-{\mathsf{a}}\|\leq\|{\mathsf{a}}\|_{\{x\},k+2n+2}\sum_{X\nsubseteq B_{r}(x)}(1+\operatorname{diam}(X)+d(x,X))^{-k-2n-2}$ (152)
@@
$\leq C(1+r)^{-k}\|{\mathsf{a}}\|_{\{x\},k+2n+2}$ (153)
where in the last line we used Lemma 8. This proves (150). To prove (151) suppose $\|{\mathsf{a}}\|^{KS}_{x,k}<\infty$ and let $X$ be any brick. Set $r:=\lfloor(\operatorname{diam}(X)+d(x,X))/4\rfloor$. Then $X\nsubseteq B_{r}(x)$. Indeed, if $X\subseteq B_{r}(x)$ then $d(x,X)\leq r$ and $\operatorname{diam}(X)\leq 2r$, implying $\operatorname{diam}(X)+d(x,X)\leq 3r$, which is impossible. Choose ${\mathsf{b}}\in{{\mathfrak{d}}_{l}}(B_{r}(x))$ with $\|{\mathsf{a}}-{\mathsf{b}}\|\leq(1+r)^{-k}\|{\mathsf{a}}\|_{x,k}^{KS}$. Since $X\nsubseteq B_{r}(x)$ we have ${\mathsf{b}}^{X}=0$, and so
@@
$\leq 4^{n+k}\|{\mathsf{a}}\|_{x,k}^{KS}(1+\operatorname{diam}(X)+d(x,X))^{-k},$ (154)
where in the second line we used [19, Proposition C.1]. ∎
A.2 Proof of Proposition 49
We begin by describing the construction of $\mathcal{J}$ and $\mathcal{K}$. Suppose $\psi$ is gapped with Hamiltonian $\mathsf{H}$} and gap $\Delta$, and write $\tau_{t}$ for the one-parameter family of LGAs obtained by exponentiating $\mathsf{H}$}. There exists [^14] See for instance Lemma 2.3 in [4]. a function $w_{\Delta}:{\mathbb{R}}\to{\mathbb{R}}$ such that $w_{\Delta}(t)=O(|t|^{-\infty})$, and the Fourier transform [^15] We use the convention $\widehat{f}(\xi)=\int f(t)e^{-i\xi t}dt$. $\widehat{w_{\Delta}}$ is supported in the interval $\Delta/2,\Delta/2]$ and satisfies $\widehat{w_{\Delta}}(0)=1$. Define $W_{\Delta}(t)$ for $t>0$ as $W_{\Delta}(t)=-\int_{t}^{\infty}w_{\Delta}(u)du$ and extend to the whole real line as an odd function. Then we define $\mathcal{J}$ and $\mathcal{K}$ as the following integral transforms:
$\mathcal{J}({\mathsf{F}}):=\int w_{\Delta}(t)\tau_{t}({\mathsf{F}})dt,$ (155)
$\mathcal{K}({\mathsf{F}}):=\int W_{\Delta}(t)\tau_{t}({\mathsf{F}})dt.$ (156)
Proof of Proposition 49.
$\mathcal{J}$ and $\mathcal{K}$ correspond to $\mathscr{J}_{{\mathsf{H}},w_{\Delta}}$ and $\mathscr{J}_{{\mathsf{H}},W_{\Delta}}$ in Section 4.1 of [19]. Part $i)$ follows from the definition of $\mathcal{K}$ and the fact that $\mathsf{H}$} preserves $\psi$. Part ii)ii) follows from Proposition 119 and Lemma 121, together with [19, Lemma F.1] (specifically line (177) therein). Part $iii)$ is [19, line (72)]. ∎
@@ Appendix B Inhomogeneous Maurer-Cartan equation
Let $({\mathfrak{M}},{\mathsf{B}})$ be a pointed pronilpotent DGLA which is a limit of an inverse system of nilpotent pointed DGLAs $({\mathfrak{M}}_{N},{\mathsf{B}}_{N})$, $N\in{\mathbb{N}}$. The set ${\rm MC}({\mathfrak{M}},{\mathsf{B}})$ of $\mathsf{B}$}-twisted MC elements has an additional equivalence relation. This section revolves around this additional structure leading eventually to a proof of Prop. 105.
@@
A $\mathsf{B}$}-twisted MC element $\mathsf{p}$} gives rise to a degree 1-1 derivation $\partial_{\mathsf{p}}=\partial+{\rm ad}_{\mathsf{p}}$ of $\mathfrak{M}$} which squares to zero (twisted differential).
Lemma 122.
@@
We are going to define an equivalence relation on ${\rm MC}({\mathfrak{M}},{\mathsf{B}})$ and ${\rm MC}({\mathfrak{M}}_{N},{\mathsf{B}}_{N})$ for all $N$. This is done in the same way as for the ordinary (homogeneous) Maurer-Cartan equation [15, 25].
@@
Second, there is a Lie algebra homomorphism from ${\mathfrak{M}}_{0}$ (resp. ${\mathfrak{M}}_{N,0}$) to the Lie algebras of affine-linear vector fields on ${\mathfrak{M}}_{-1}$ (resp. ${\mathfrak{M}}_{N,-1}$). This homomorphism maps ${\mathsf{a}}\in{\mathfrak{M}}_{0}$ or ${\mathfrak{M}}_{N,0}$ to the affine-linear vector field
$\xi_{\mathsf{a}}({\mathsf{p}})=[{\mathsf{a}},{\mathsf{p}}]-\partial{\mathsf{a}},$ (157)
@@
We say that elements ${\mathsf{p}}_{1},{\mathsf{p}}_{2}$ of ${\rm MC}({\mathfrak{M}},{\mathsf{B}})$ or ${\rm MC}({\mathfrak{M}}_{N},{\mathsf{B}}_{N})$ are equivalent if they belong to the same orbit of these actions.
Remark 124.
@@
Now come the interesting statements. Assume from now on that the DGLAs ${\mathfrak{M}}_{N}$ and $\mathfrak{M}$} are acyclic, that ${\mathfrak{j}}_{N}\subseteq{\mathfrak{M}}_{N}$ is central for all $N>1$, and that the morphisms $r_{N,N-1}$ are surjective for all $N>1$.
Lemma 126.
@@
The only if statement follows from Lemma 122. To prove the if direction, we use induction on $N$. Assume ${\rm MC}({\mathfrak{M}}_{N-1},{\mathsf{B}}_{N-1})$ is non-empty. Let ${\mathsf{p}}_{N-1}\in{\rm MC}({\mathfrak{M}}_{N-1},{\mathsf{B}}_{N-1})$. Pick $\tilde{\mathsf{p}}_{N}\in r_{N,N-1}^{-1}({\mathsf{p}}_{N-1})$. Since ${\mathsf{p}}_{N-1}$ is a ${\mathsf{B}}_{N-1}$-twisted MC element and $r_{N,N-1}({\mathsf{B}}_{N})={\mathsf{B}}_{N-1}$, we must have
$\partial\tilde{\mathsf{p}}_{N}+\frac{1}{2}[\tilde{\mathsf{p}}_{N},\tilde{\mathsf{p}}_{N}]={\mathsf{B}}_{N}+{\mathsf{q}}_{N}$ (159)
for some ${\mathsf{q}}_{N}\in{\mathfrak{j}}_{N}$. We look for a solution of the ${\mathsf{B}}_{N}$-twisted MC equation of the form ${\mathsf{p}}_{N}=\tilde{\mathsf{p}}_{N}+{\mathsf{b}}_{N}$ where ${\mathsf{b}}_{N}\in{\mathfrak{j}}_{N}$. Taking into account that ${\mathfrak{j}}_{N}$ is central, the inhomogeneous MC equation reduces to $\partial{\mathsf{b}}_{N}=-{\mathsf{q}}_{N}$. Since by assumption ${\mathfrak{j}}_{N},{\mathfrak{M}}_{N}$, and ${\mathfrak{M}}_{N-1}$ form a short exact sequence and the latter two are acyclic, so is ${\mathfrak{j}}_{N}$. Hence such a ${\mathsf{b}}_{N}$ exists. This completes the inductive step proving that ${\rm MC}({\mathfrak{M}}_{N},{\mathsf{B}}_{N})\neq\varnothing$ for all $N$. Moreover, we also proved that the morphisms ${\rm MC}({\mathfrak{M}}_{N},{\mathsf{B}}_{N})\rightarrow{\rm MC}({\mathfrak{M}}_{N-1},{\mathsf{B}}_{N-1})$ are surjective. Therefore by Lemma 122 ${\rm MC}({\mathfrak{M}},{\mathsf{B}})$ is non-empty. ∎
The above lemma proves the first part of Prop. 105. Indeed when ${\mathfrak{M}}_{1},{\mathfrak{M}}_{1}]=0$, ${\rm MC}({\mathfrak{M}}_{1},{\mathsf{B}}_{1})\neq\varnothing$ is implied by acyclicity. Next we show that with the above assumption on ${\mathfrak{M}}_{N}$ and $\mathfrak{M}$} all $\mathsf{B}$}-twisted MC elements are equivalent.
Lemma 127.
@@
$\exp({\mathsf{b}}+{\mathsf{a}})*({\mathsf{p}})=\exp({\mathsf{b}})*({\mathsf{p}})-\partial{\mathsf{a}}.$ (160)